In triangle ABC, C is the right angle, B = 30 degrees, and side b = 3. Find the exact value of the missing sides. Thank you for your help.
Answer by venugopalramana(3286)

In triangle ABC, C is the right angle, B = 30 degrees, and side b = 3. Find the exact value of the missing sides. Thank you for your help.
WE HAVE SIN(B)=SIN(30)=0.5=AC/AB=b/c=3/c
AB = c =3/0.5=6
TAN(B)=TAN(30)=SQRT(3)=1.7321=AC/BC=b/a=3/a
BC = a = 3/1.7321=1.7321